# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v5

@@ Abstract
Large Multimodal Models (LMMs) have shown significant visual reasoning capabilities by connecting a visual encoder and a large language model. LMMs typically take in a fixed and large amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content. Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which further increases the number of visual tokens significantly. However, due to the inherent design of the Transformer architecture, the computational costs of these models tend to increase quadratically with the number of input tokens. To tackle this problem, we explore a token reduction mechanism that identifies significant spatial redundancy among visual tokens. In response, we propose PruMerge, a novel adaptive visual token reduction strategy that significantly reduces the number of visual tokens without compromising the performance of LMMs. Specifically, to metric the importance of each token, we exploit the sparsity observed in the visual encoder, characterized by the sparse distribution of attention scores between the class token and visual tokens. This sparsity enables us to dynamically select the most crucial visual tokens to retain. Subsequently, we cluster the selected (unpruned) tokens based on their key similarity and merge them with the unpruned tokens, effectively supplementing and enhancing their informational content. Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 14 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
@@ 1 Introduction
Large Language Models (LLMs) [OpenAI, 2023b, Team et al., 2023, Jiang et al., 2023, Touvron et al., 2023] have shown strong reasoning abilities. LLMs are usually high-capacity Transformers [Vaswani et al., 2017] pretrained with a large-scale text corpus. Large Multimodal Models (LMMs), inherit LLMs for text generation, while also leveraging a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
LMMs need substantial computation for inference. The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM. For example, the commonly used CLIP visual encoder, ViT-L, only has 0.3B parameters, while the corresponding LLM such as LLaMA [Touvron et al., 2023] or Vicuna [Vicuna, 2023] can have 7B or 13B parameters. As a result, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
Prior works [Chu et al., 2023, 2024, Yuan et al., 2023a] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [Javaheripi et al., 2023]. However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap on visual question-answering and reasoning tasks such as VQAv2 and MM-Bench [Chu et al., 2024]. A similar approach is to apply quantization for LLMs [Liu et al., 2023b, Yuan et al., 2024].
However, the cost of LLMs comes from not only its large number of parameters, but also the length of the input context due to the quadratic complexity of the Transformer’s attention operation. The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens. For example, in LLaVA-1.5, 576 visual tokens are appended, and in Video-LLaVA [Lin et al., 2023] that number is even higher, leading to high training and inference costs. Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
(a) Main idea of PurMerge.
(b) PruMerged Token Visualization.
@@
Figure 1: (a) We prune and merge visual tokens produced by the vision encoder, while keeping all other procedures of the LMM the same. By reducing the number of visual tokens, PruMerge, significantly reduces the computation cost for text generation in LMMs (around 4-10 times in FLOPs for LMM prefill), while can maintain comparable performance. (b) A visualization of the selected tokens. PruMerge can adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently. More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images. The attentive tokens are usually located at the regions with dense information.
@@
In our study, we find that many visual tokens are redundant, similar to findings in previous related work [Bolya et al., 2023, Liu et al., 2022], and most of the visual tokens can be pruned with little sacrifice in performance. In particular, the similarity (i.e., attention scores in the visual encoder’s self-attention module) between the class token and spatial patches are sparse, indicating that only a small number of visual tokens are related to key visual information in the visual samples. Motivated by this, we use this sparse similarity to adaptively select important visual tokens, as shown in Fig.1(b). Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune unimportant visual tokens. Moreover, we merge the visual tokens using $k$-nearest neighbor and update the selected important visual tokens via weighted averaging, which further enhances performance. Finally, we design PruMerge+, which samples visual tokens spatial-uniformly to complement the unpruned tokens. PruMerge+ not only minimizes performance degradation but also ensures substantial token reduction, maintaining a more comprehensive and representative selection of visual tokens.
@@
Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens for an image on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Furthermore, PruMerge showcases its versatility across various modalities, including video. By integrating PruMerge with Video-LLaVA during the inference phase alone—eliminating the need for additional training—we not only expedite processing within video-LLMs but also enhance their performance across multiple benchmarks.
@@ 2.1 Efficient Large Multimodal Models (LMMs)
Large Language Models (LLMs) such as GPT-4 [OpenAI, 2023b], LLaMA [Touvron et al., 2023], Mistral [Jiang et al., 2023], and Gemini [Team et al., 2023] have demonstrated strong question answering and reasoning capabilities over text. Large Multimodal Models (LMMs) [Liu et al., 2023b, Zhu et al., 2023, Yin et al., 2023, Zhang et al., 2024] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format. More recent works extend whole-image understanding into region-level understanding [Cai et al., 2024, Zhang et al., 2023b, Peng et al., 2023, Chen et al., 2023], video understanding [Lin et al., 2023, Zhang et al., 2023a] and 3D scene understanding [Hong et al., 2023]. Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either an MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022]. The number of visual tokens can be prohibitively long, especially when the images are high-resolution [Liu et al., 2024, OpenAI, 2023a]. In this paper, we reduce the number of visual tokens with a novel adaptive prune and merge procedure.
While LMMs have made significant advances, their large-scale training and deployment incur significant computational costs, requiring efficient parallel device implementations. Google’s Gemini [Team et al., 2023] is a pioneer in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices, although it is not open-source. Open-source alternatives like LLaVA-1.5 [Liu et al., 2023a] employ advanced compression techniques such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024]. MobileVLM [Chu et al., 2023] and its improved version, MobileVLM-v2 [Chu et al., 2024], focus on compact architecture designs and training optimizations for mobile use. Similarly, TinyGPT-V [Yuan et al., 2023a] uses the Phi-2 [Javaheripi et al., 2023] LLM backbone to exceed the performance of larger models. LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for better generalizability. TinyLLaVA [Zhou et al., 2024] explores the effects of architectural choices and training optimizations, achieving comparable performance to larger models. MoE-LLaVA [Lin et al., 2024] incorporates a Mixture of Experts to address model sparsity, enhancing efficiency and performance.
@@ 2.2 Token Reduction
The notorious quadratic complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the input sequence length. Sparse attention methods such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context. Token merging [Bolya et al., 2023] utilizes full attention but gradually reduces the number of tokens in each transformer block by selecting the most representative tokens with bipartite matching. One of the main causes of the inefficiency of recent LMMs is their use of a large number of prefix visual tokens that serve as a fixed budget for context [Liu et al., 2023b, Zhu et al., 2023]. In our study, we present a novel plug-and-play token reduction method based on visual token similarities, which achieves comparable performance with less than one tenth of the original tokens.
@@ 3.1 Preliminaries
Figure 2: The conceptual idea of PruMerge. Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens via an outlier detection algorithm (see Sec.3.2); (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster (see Sec.3.3). Here $m$ denotes the visual token compression ratio.
Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for LMMs, in which the input image is converted into a sequence of representative tokens by the ViT, and then fed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024]. An input image is divided into a grid of patches and each patch is projected into a token embedding by the ViT. In addition to the patch tokens, a class token (i.e., [CLS] token) is computed to aggregate global image information for classification. A ViT consists of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), skip connections, and layer normalization [Ba et al., 2016]. These components work together to improve the model’s capability to understand visual data [Han et al., 2022]. In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, respectively. The self-attention computes the relevance of each item to other items:
$\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (3.1)
@@
$\mathbf{a}_{\texttt{cls}}=\text{softmax}\left(\frac{\mathbf{q}_{\texttt{cls}}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right).$ (3.2)
@@
$\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3.3)
@@
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which maps the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024]. The overall architecture of an LMM is depicted in Figure 1.
It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022]. Mathematically, if there are $N$ tokens in the input, the self-attention mechanism computes a $N\times N$ matrix of attention scores, where each entry in this matrix represents the attention score between a pair of tokens. However, there is an increasing demand for processing high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
@@ 3.2 Adaptive Important Token Selection via Outlier Detection
Figure 3: (a) The distribution of attention scores (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens. The values in the y-axis are logarithmic values. Notably, most of the spatial visual tokens have near-zero attention values with the class token. (b) Visualizations of PruMerge and PruMerge+.
The most straightforward solution to improve the efficiency of visual token utilization in LMMs is to prune redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022]. To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
As discussed in Sec. 3.1, LMMs typically leverage an extensive stack of visual tokens to represent the visual information. On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [Radford et al., 2021] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token. To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 3.2. Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3(a). This sparse distribution underpins our methodology for identifying crucial visual tokens. By employing outlier detection algorithms, we aim to adaptively select visual tokens that best represent an image’s features effectively.
Interquartile Range (IQR) Method for outlier detection. To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [Boukerche et al., 2020], a statistical technique known for its robustness in outlier detection. The essence of the IQR method lies in its capability to establish a boundary or “fence” that delineates the normal range of data. This is achieved by calculating the IQR (the difference between the third quartile Q3 and the first quartile Q1) and subsequently defining the outer limits of the normal range as 1.5 times the IQR above Q3 and below Q1. Specifically, the computation is as follows: the “lower fence” is set at 1.5 $\times$ IQR below Q1, and the “upper fence” is set at 1.5 $\times$ IQR above Q3. Any attention values residing outside these fences are classified as outliers. In practice, only the “upper fence” is activated since the attention score is positive. Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context. Note that we use the class attention value from the penultimate layer for this calculation.
As shown in Figure 1(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image. Simpler images such as “Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens. (2) The sampled tokens are typically spatially aligned with important content. Such visualizations align with our visual token sampling design. These trends are also observed at the benchmark level; in Table 6, the average token numbers on various benchmarks differ.

### v6

@@ Abstract
Large Multimodal Models (LMMs) have shown significant visual reasoning capabilities by connecting a visual encoder and a large language model. LMMs typically take in a fixed and large amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content. Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which further increases the number of visual tokens significantly. However, due to the inherent design of the Transformer architecture, the computational costs of these models tend to increase quadratically with the number of input tokens. To tackle this problem, we explore a token reduction mechanism that identifies significant spatial redundancy among visual tokens. In response, we propose PruMerge, a novel adaptive visual token reduction strategy that significantly reduces the number of visual tokens without compromising the performance of LMMs. Specifically, to metric the importance of each token, we exploit the sparsity observed in the visual encoder, characterized by the sparse distribution of attention scores between the class token and visual tokens. This sparsity enables us to dynamically select the most crucial visual tokens to retain. Subsequently, we cluster the selected (unpruned) tokens based on their key similarity and merge them with the unpruned tokens, effectively supplementing and enhancing their informational content. Empirically, when applied to LLaVA-1.5 [28] and Video-LLaVA [26], our approach can reduce the number of visual tokens by 4 times, and achieve comparable or better performance across diverse visual question-answering and reasoning tasks.
@@ 1 Introduction
Large Language Models (LLMs) [34, 42, 21, 43] have shown strong reasoning abilities. LLMs are usually high-capacity Transformers [44] pretrained with a large-scale text corpus. Large Multimodal Models (LMMs), inherit LLMs for text generation, while also leveraging a visual encoder such as CLIP-ViT [37] to embed image patches into visual tokens as the prefix visual context.
@@
Figure 1: Main idea of PruMerge: We prune and merge visual tokens produced by the vision encoder, while keeping all other procedures of the LMM the same. By reducing the number of visual tokens, PruMerge, significantly reduces the computation cost for text generation in LMMs (around 4-10 times in FLOPs for LMM prefill), while can maintain comparable performance. A visualization of the selected tokens: PruMerge can adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently. More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images. Attentive tokens are located at the regions with dense information.
LMMs need substantial computation for inference. The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM. For example, the commonly used CLIP visual encoder only has 0.3B parameters, while the corresponding LLM such as LLaMA [43], Vicuna [45], DeepSeek [27], and Qwen [3] can have 7B, 13B, or even 70B+ parameters. Therefore, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
Prior works [8, 9, 49] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [20]. However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap on visual question-answering and reasoning tasks such as VQAv2 and MM-Bench [9]. A similar approach is to apply quantization for LLMs [30, 51].
However, the cost of LLMs comes from not only its large number of parameters, but also the length of the input context due to the quadratic complexity of the Transformer’s attention operation. The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens. For example, in LLaVA-1.5, 576 visual tokens are appended, and in Video-LLaVA [26] that number is even higher (2048), leading to high training and inference costs. Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
@@
In our study, we find that many visual tokens are redundant, similar to findings in previous related work [4, 31], and most of the visual tokens can be pruned with little sacrifice in performance. In particular, the similarity (i.e., attention scores in the visual encoder’s self-attention module) between the class token and spatial patches are sparse, indicating that only a small number of visual tokens are related to key visual information in the visual samples. Motivated by this, we use this sparse similarity to adaptively select important visual tokens, as shown in Fig.1. Specifically, we leverage the Interquartile Range (IQR) [5] scoring function in outlier detection to prune unimportant visual tokens. Moreover, we merge the visual tokens using $k$-nearest neighbor and update the selected important visual tokens via weighted averaging, which further enhances performance. Finally, we design PruMerge+, which samples visual tokens spatial-uniformly to complement the unpruned tokens. PruMerge+ not only minimizes performance degradation but also ensures substantial token reduction, maintaining a more comprehensive and representative selection of visual tokens.
@@
Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [28], where with just 5.5% of visual tokens, which is around 32 tokens for an image on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Furthermore, PruMerge showcases its versatility across various modalities, including video. By integrating PruMerge with Video-LLaVA [26] during the inference phase alone (i.e., no need for additional training) we not only expedite processing within video-LLMs but also enhance their performance across multiple benchmarks.
@@ 2.1 Efficient Large Multimodal Models (LMMs)
Large Language Models (LLMs) such as GPT-4 [34], LLaMA [43], Mistral [21], and Gemini [42] have demonstrated strong question answering and reasoning capabilities over text. Large Multimodal Models (LMMs) [30, 55, 48, 52] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format. More recent works extend whole-image understanding into region-level understanding [6, 54, 36, 7], video understanding [26, 53] and 3D scene understanding [18]. Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either an MLP [28], Qformer [10, 55], or resampler [1]. The number of visual tokens can be prohibitively long, especially when the images are high-resolution [29, 35]. In this paper, we reduce the number of visual tokens with a novel adaptive prune and merge procedure.
While LMMs have made significant advances, their large-scale training and deployment incur significant computational costs, requiring efficient parallel device implementations. Google’s Gemini [42] is a pioneer in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices, although it is not open-source. Open-source alternatives like LLaVA-1.5 [28] employ advanced compression techniques such as 4/8 bit quantization [11, 38]. MobileVLM [8] and its improved version, MobileVLM-v2 [9], focus on compact architecture designs and training optimizations for mobile use.
@@ 2.2 Token Reduction
The quadratic complexity of Transformers [44] poses a significant challenge in scaling input sequence length. Various approaches try to address this issue. Sparse attention methods, e.g., Linformer [46] and ReFormer [22], reduce complexity by limiting attention operations to specific regions rather than the full context. Token reduction can also accelerate Transformers [17]. Methods like [31, 47, 25, 4, 13] focus on reducing the number of tokens within the internal transformer structure, thereby decreasing computational load. For instance, token merging [4] employs full attention but progressively reduces tokens in each transformer block by selecting the most representative tokens through bipartite matching. However, these uni-modal token reduction methods are not directly applicable to LMMs. One of the main inefficiencies in LMMs stems from their use of numerous prefix visual tokens as a fixed context budget [30, 55] (analyzed further in Sec. 4.2), not from the internal structure of Transformers. We discuss the unsuitability of existing uni-modal token reduction methods for LMM acceleration in Sec. 3.5. In our study, we introduce a plug-and-play token reduction method specifically designed for LMMs. Our approach, based on visual token similarities, achieves comparable performance while using less than one-fourth of the original tokens. The core of our method is a sparsity-based selection mechanism that identifies “anchor” tokens via sparse attention scores within the modality encoder, and is the most crucial design element of PruMerge. In parallel to our work, [39] proposes CrossGet, a graph-matching-based algorithm for token matching. While both approaches aim to reduce tokens in multimodal contexts, they differ significantly in their methodologies. Beyond the token selection module, our token merging module also differs from CrossGet’s graph soft matching. Our k-nearest neighbors clustering approach has a time complexity of $O(n)$, which is more computationally efficient compared to CrossGet’s $O(n^{2})$ complexity, thus enhancing scalability and efficiency.
@@ 3.1 Preliminaries
Vision Transformers (ViTs) [12] are the most widely used vision encoder for LMMs, in which the input image is converted into a sequence of representative tokens by the ViT, and then fed into an LLM for understanding [29, 55, 18, 52]. An input image is divided into a grid of patches and each patch is projected into a token embedding by the ViT. In addition to the patch tokens, a class token (i.e., [CLS] token) is computed to aggregate global image information for classification. A ViT consists of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), skip connections, and layer normalization [2]. These components work together to improve the model’s capability to understand visual data [16]. In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, respectively. The self-attention computes the relevance of each item to other items:
$\mathbf{Y}=\text{Self-Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathbf{A}\cdot\mathbf{V}$ (1)
@@
$\mathbf{a}_{\texttt{cls}}=\text{softmax}\left(\frac{\mathbf{q}_{\texttt{cls}}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right).$ (2)
@@
$\text{FFN}(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X})$ (3)
@@
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which maps the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [52]. The overall architecture of an LMM is depicted in Figure 1.
Importantly, the computational cost with these models increases quadratically with the number of input tokens to the LLM [41]. Mathematically, if there are $N$ tokens in the input, the self-attention mechanism computes a $N\times N$ matrix of attention scores, where each entry in this matrix represents the attention score between a pair of tokens. However, there is an increasing demand for processing high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
Figure 2: PruMerge has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens via an outlier detection algorithm (see Sec.3.2); (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster (see Sec.3.3). Here $m$ denotes the visual token compression ratio.
@@ 3.2 Adaptive Important Token Selection via Outlier Detection
The most straightforward solution to improve the efficiency of visual token utilization in LMMs is to prune redundant visual tokens [31, 47, 25]. To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
As discussed in Sec. 3.1, LMMs typically leverage an extensive stack of visual tokens to represent the visual information. On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [37] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token. To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 2. Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3(a). This sparse distribution underpins our methodology for identifying crucial visual tokens. By employing outlier detection algorithms, we aim to adaptively select visual tokens that best represent an image’s features effectively.
Interquartile Range (IQR) Method for outlier detection. To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [5], a statistical technique known for its robustness in outlier detection. Its essence lies in its capability to establish a boundary or “fence” that delineates the normal range of data. This is achieved by calculating the IQR (the difference between the third quartile Q3 and the first quartile Q1) and subsequently defining the outer limits of the normal range as 1.5 times the IQR above Q3 and below Q1. Specifically, the computation is as follows: the “lower fence” is set at 1.5 $\times$ IQR below Q1, and the “upper fence” is set at 1.5 $\times$ IQR above Q3. Any attention values residing outside these fences are classified as outliers. In practice, only the “upper fence” is activated since the attention score is positive. Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context. Note that we use the class attention value from the penultimate layer for this calculation.
As shown in Fig.3(b), the sampled visual tokens show two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image. Simpler images such as “Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens. (2) The sampled tokens are typically spatially aligned with important content. Such visualizations align with our visual token sampling design. These trends are also observed at the benchmark level; in Tab.6, the average token numbers on various benchmarks differ.
@@
Figure 3: (a) Distribution of attention scores (in CLIP-ViT’s penultimate layer) between the [cls] token and visual tokens. The y-axis shows logarithmic values. Notably, most spatial visual tokens have near-zero attention values with the class token. (b) Visualizations of PruMerge and PruMerge+.